OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. x2 - 16 = 0 and x2 = -2x + 24 Part 2: Describe what the solution(s) represent to the graph of each. Part 3: How are the graphs alike? How are they different?

OpenStudy (anonymous):

it is good

OpenStudy (anonymous):

Okay but I am not sure how to do the rest :(

OpenStudy (anonymous):

b) the solutions represent the x-intercepts or the zeros

OpenStudy (anonymous):

with the another equation you can do this

OpenStudy (anonymous):

x2 = -2x + 24 => x2+2x-24=0 from here you could factor or use the quadratic formula

OpenStudy (anonymous):

So is that x2 = -2x + 24 then? I need that answered :) I am not 100% sure how to do this

OpenStudy (anonymous):

yes x2 = -2x + 24 is the same that x2+2x-24=0 from here you can factor and get (x+6)(x-4) so the solutions are x=-6 and x=4

OpenStudy (anonymous):

it is the same answer for the b part

OpenStudy (anonymous):

now for the c) part you can use this

OpenStudy (anonymous):

http://graph.seriesmathstudy.com/

OpenStudy (anonymous):

in this page u can make the graph and see how the graphs alike or differ

OpenStudy (anonymous):

I don't know how to use this :( sorry.

OpenStudy (anonymous):

ok i will atach the file

OpenStudy (anonymous):

the red one is x2+2x-24

OpenStudy (anonymous):

well the vertex (the minimum point on the graph) are diferent and one of the zeros are diferent

OpenStudy (anonymous):

i will leave take kare

OpenStudy (anonymous):

Thanks so much!!!!